Stephanie Corporation sells a single product. Budgeted sales for the year are anticipated to be 669,000 units, estimated beginning inventory is 104,000 units, and desired ending inventory is 87,000 units. The quantities of direct materials expected to be used for each unit of finished product are given below.

Material A: 0.50 lb. per unit @ $0.72 per pound
Material B: 1.00 lb. per unit @ $1.61 per pound
Material C: 1.20 lb. per unit @ $1.08 per pound

The dollar amount of Material A used in production during the year is

a.$844,992

b.$1,049,720

c.$234,720

d.$240,840
Answer #1

Units to be produced

= Budgeted sales + Desired ending inventory - Beginning inventory

= 669,000+87,000-104,000 = 652,000

Raw materials needed = 652,000*0.50 = 326,000

In dollars = 326,000*0.72

= 234,720